OpenStudy (anonymous):

What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms?

OpenStudy (anonymous):

Ooh well let's get the sequence first. What do you see is the pattern in here?

OpenStudy (anonymous):

Times 2

OpenStudy (anonymous):

no times -2

OpenStudy (anonymous):

Close. It's not quite 2. How would we get from 8 to -16 and then back to 32?

OpenStudy (anonymous):

Oh you got it

OpenStudy (anonymous):

So now the only way I know how to do it is find multiply it all be -2

OpenStudy (anonymous):

so that is 8,-16,32,-64,128,-256,512,-1024,2048,-4096,8192,-16384,32768,-65536,131072

OpenStudy (cwrw238):

the formula for the sum of n terms is Sn = a * 1 - r^n ------ 1 - r

OpenStudy (anonymous):

I knew there was an easier way :) Thank you

OpenStudy (cwrw238):

a = first term, r = common ratio ( here = -2)

OpenStudy (anonymous):

ok im going to try and work it out

OpenStudy (anonymous):

Sn= a*1 - r^n ------------- 1- r Sn= 8*1 - (-)2^n ------------- 1- (-)2 Sn= 8 + 2^n ------------- 3

OpenStudy (anonymous):

thats what i got........ ?

ganeshie8 (ganeshie8):

Sn= a*1 - r^n ------------- 1- r Sn= 8*1 - (-2)^n ------------- 1- (-2)

ganeshie8 (ganeshie8):

since you wanto find sum for 15 terms, put n = 15

OpenStudy (anonymous):

ohh okk . one sec.

OpenStudy (anonymous):

Sn= 8 + 2^15 ------------- 3 Sn= 8 + 30 ------------- 3 Sn= 38 ------------- 3

ganeshie8 (ganeshie8):

not exactly...

ganeshie8 (ganeshie8):

2^15 means, 2 power 15

ganeshie8 (ganeshie8):

like this : \(2^{15}\)

ganeshie8 (ganeshie8):

Sn= a*1 - r^n ------------- 1- r Sn= 8*1 - (-2)^n ------------- 1- (-2) Sn= 8*1 - (-2)^15 ------------- 1- (-2) Sn= 8*1 + (2)^15 ------------- 1 + (2) Sn= 8*1 + 32768 ------------- 3 Sn= 8*32769 ------------- 3

ganeshie8 (ganeshie8):

see if u can take it from here

OpenStudy (cwrw238):

thats * not + maybe my formula was ambiguous i should have written it as Sn = a * (1 - r^n) ------- ( 1 - r)

OpenStudy (cwrw238):

so its 8 * 32769 -------- 3